OpenStudy (anonymous):

The variable complex number ß is given by ß=1+cos 2 θ +isin2 θ , where θ takes all values in the interval−1 2 π < θ <1 2 π . (i) Show that the modulus of ß is 2 cos θ and the argument of ß is θ

OpenStudy (michele_laino):

is beta like this: \[\Large \beta = 1 + \cos \left( {2\theta } \right) + i\sin \left( {2\theta } \right)\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! Then by definition, the modulus of \beta is given by the subsequent formula: \[\Large \left| \beta \right| = \sqrt {{{\left\{ {1 + \cos \left( {2\theta } \right)} \right\}}^2} + {{\left\{ {\sin \left( {2\theta } \right)} \right\}}^2}} = ...?\]

OpenStudy (anonymous):

i am not being able to solve... help pls ?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

here is your steps: \[ \begin{gathered} {\left\{ {1 + \cos \left( {2\theta } \right)} \right\}^2} + {\left\{ {\sin \left( {2\theta } \right)} \right\}^2} = 1 + {\left( {\cos \left( {2\theta } \right)} \right)^2} + 2\cos \left( {2\theta } \right) + {\left( {\sin \left( {2\theta } \right)} \right)^2} = \hfill \\ = 2 + 2\cos \left( {2\theta } \right) = 2\left( {1 + \cos \left( {2\theta } \right)} \right) = 2 \times 2{\left( {\cos \theta } \right)^2} = 4{\left( {\cos \theta } \right)^2} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so we can write: \[\Large \left| \beta \right| = \sqrt {4{{\left( {\cos \theta } \right)}^2}} = 2\left| {\cos \theta } \right|\]

OpenStudy (anonymous):

thank you :) and the second part ?

OpenStudy (michele_laino):

for second part, we can write, I call with \alpha the principal argument of \beta, so I can write: \[\Large \begin{gathered} \cos \alpha = \frac{{1 + \cos \left( {2\theta } \right)}}{{2\cos \theta }}, \hfill \\ \\ \sin \alpha = \frac{{\sin \left( {2\theta } \right)}}{{2\cos \theta }} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

then we get: \[\Large \tan \alpha = \frac{{\sin \left( {2\theta } \right)}}{{2\cos \theta }} \times \frac{{2\cos \theta }}{{1 + \cos \left( {2\theta } \right)}} = \frac{{\sin \left( {2\theta } \right)}}{{1 + \cos \left( {2\theta } \right)}}\]

OpenStudy (michele_laino):

and finally: \[\Large \begin{gathered} \tan \alpha = \frac{{\sin \left( {2\theta } \right)}}{{1 + \cos \left( {2\theta } \right)}} = \frac{{2\sin \theta \cos \theta }}{{1 + {{\left( {\cos \theta } \right)}^2} - {{\left( {\sin \theta } \right)}^2}}} = \hfill \\ \hfill \\ = \frac{{2\sin \theta \cos \theta }}{{2{{\left( {\cos \theta } \right)}^2}}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

Prove that the real part of 1 ß is constant.

OpenStudy (michele_laino):

the real part of (1+\beta) ?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

we have: \[\Large \begin{gathered} \Re \left( {1 + \beta } \right) = \frac{{1 + \beta + 1 + {\beta ^*}}}{2} = \frac{{2 + \beta + {\beta ^*}}}{2} = \hfill \\ \hfill \\ = 1 + \Re \beta \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

and: \[\Large \begin{gathered} \Re \beta = \hfill \\ = \frac{{1 + \cos \left( {2\theta } \right) + i\sin \left( {2\theta } \right) + 1 + \cos \left( {2\theta } \right) - i\sin \left( {2\theta } \right)}}{2} = \hfill \\ \hfill \\ = 1 + \cos \left( {2\theta } \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so: \[\Large \Re \left( {1 + \beta } \right) = 1 + \Re \beta = 2 + \cos \left( {2\theta } \right)\]

OpenStudy (michele_laino):

which is not constant, since it depends on \theta